# Analytic Continual Test-Time Adaptation for Multi-Modality Corruption

### v1

@@ Appendix B Pseudo-code for MDAA
In this appendix, we provide the pseudo-code for our MDAA in Algorithm 1. For the pre-trained model, we integrate an individual Analytic Classifier (AC) for each network block, using the source dataset to initialize the classifiers as well as the memory bank. During the inference and adaptation periods, the model reconstructs the output labels for each sample using the Soft Pseudo-label Strategy (SPS) and determines which ACs need to be updated through the Dynamic Selecting Mechanism (DSM).
Appendix C Benchmarks,Backbone and Implementation Details
C.1 Details about the Benchmarks
@@
ALL experiments are conducted on the two popular multi-modal datasets Kinetics (Kay et al., 2017) and VGGSound (Chen et al., 2020). Yang et al. (2024) further provides their corrupted visual and audio modality for TTA tasks.
@@
Both datasets’ visual and audio modalities were extracted following the method described in Gong et al. (2022b). To systematically explore the distributional shifts across modalities, various corruption types were applied to both visual and audio components. Following Yang et al. (2024), 15 corruption types were introduced for the visual modality, each with five levels of severity for comprehensive evaluation. These corruptions include ”gaussian noise”, ”shot noise”, ”impulse noise”, ”defocus blur”, ”glass blur”, ”motion blur”, ”zoom blur”, ”snow”, ”frost”, ”fog”, ”brightness”, ”contrast”, ”elastic transform”, ”pixelate”, and ”jpeg compression”. Similarly, the audio modality was subjected to 6 different corruptions: ”gaussian noise”, ”traffic noise”, ”crowd noise”, ”rain”, ”thunder”, and ”wind”. The corrupted versions of these benchmarks are referred to as Kinetics50-C and VGGSound-C, respectively. Visualizations of sample corrupted video frames and audio spectrograms are provided in Fig.5 and Fig.6.
C.2 CAV-MAE Backbone
@@
CAV-MAE is employed as the pre-trained model for MM-CTTA in this paper. Its architecture consists of 11 Transformer blocks (known as feature encoder networks) dedicated to modality-specific feature extraction, followed by an additional Transformer block (known as fusion network) responsible for cross modal fusion. For the video input, 10 frames are sampled from each clip, from which a single frame is randomly selected and fed into the Transformer encoder for the visual modality. In the case of the audio input, the original 10-second audio waveform is transformed into a 2D spectrogram before being processed by the Transformer encoder for the audio modality (Gong et al., 2022a).
C.3 Implementation Details
@@
In the final step of Algorithm 1, we determine the hyperparameters for MDAA. The expansion layer dimension, denoted as , theoretically benefits from larger values. However, an excessively large dimension may introduce a significant number of parameters, increasing the computational load. Given the constraints of our available GPU resources, we set the dimension of to 8000. The necessity of the parameter in 3.2 has been established in Zhuang et al. (2022). The model demonstrates stable performance over a wide range of values, indicating that as long as is within a reasonable range, its impact on the model’s performance remains minimal. After conducting a sweep over five orders of magnitude (), we set to 1 for Kinetics-50 and 10 for VGGSound. As discussed in Sections 4.2 and 4.3, the threshold in DSM is fixed at 0.001 for both datasets, while the parameter in SPS is set to 2 for Kinetics-50 and 7 for VGGSound. All experiments were conducted on an RTX3090 GPU, with results averaged over three runs.
Appendix D More Experiment Results
@@
In this appendix we provide four more experiments for reference where Appendix .D.1 and D.2 are the supplementary comparative studies while Appendix .D.3 and D.4 are the further discussion on hyperparameters applied in MDAA.
D.1 Performance Comparison on Multi-modal Corruption
@@
We follow the experimental setup in (Yang et al., 2024) to examine whether our model remains superior to others when the adaptation task is not continual. In this section, the corrupted target domains are treated as independent, and the results for corrupted audio and video modalities are presented in Tables 7, 8, and 9. The results show that the SOTA performance of MDAA is not only due to its ability to combat catastrophic forgetting but also its strong capacity to handle MM-TTA tasks effectively.
@@
D.2 Performance Comparison on Single-modality Continual Corruption
@@
| Method | Type | Kinetics50-C | VGGSound-C | ||||||||||||
| CoTTA | CTTA | 73.89 | 66.84 | 68.08 | 67.53 | 71.10 | 69.33 | 69.46 | 39.70 | 34.88 | 35.54 | 33.67 | 39.42 | 33.83 | 36.17 |
| CoTTA | CTTA | 69.88 | 67.67 | 67.84 | 68.54 | 69.63 | 70.76 | 69.05 | 25.06 | 25.08 | 28.01 | 33.22 | 37.90 | 29.67 | 29.82 |
| CoTTA | CTTA | 49.17 | 46.65 | 43.69 | 61.82 | 60.00 | 63.54 | 62.34 | 50.60 | 52.13 | 53.35 | 60.25 | 50.86 | 60.43 | 58.28 | 62.61 | 55.71 |
| CoTTA | CTTA | 56.89 | 57.26 | 56.92 | 66.39 | 65.16 | 68.11 | 66.52 | 60.23 | 62.46 | 50.36 | 73.47 | 50.02 | 66.04 | 67.85 | 67.06 | 62.31 |
| CoTTA | CTTA | 53.03 | 53.08 | 53.03 | 52.92 | 52.28 | 51.84 | 51.41 | 50.83 | 50.08 | 49.66 | 49.22 | 48.85 | 48.30 | 47.62 | 47.20 | 50.62 |
| CoTTA | CTTA | 48.10 | 48.51 | 49.13 | 49.90 | 50.47 | 50.93 | 51.67 | 52.14 | 53.17 | 53.19 | 55.86 | 53.21 | 58.02 | 56.60 | 56.71 | 52.51 |
D.3 Dynamic threshold update
The threshold we used in DSM is a fixed number. In this section we attempt to update in a dynamic way during the adaptation. We define the threshold for classifier in time as
where is the learning rate, is the initial threshold and is the batch size. is calculated to reflect the gap between leader and follower . The original intention of this design is to adjust the size of the threshold according to the change of , so as to eliminate the statistical bias of different domains. However, as shown in Table 14, such attempt is not only achieve lower performance while needs more variables to be memorized. So only the fixed threshold is used in the formal method.
@@ D.3 Dynamic threshold update
| KS-video | ||
| 0.9 | 64.29 | 69.29 |
| 0.8 | 64.88 | 69.29 |
| 0.7 | 65.43 | 69.30 |
| 0.6 | 65.20 | 68.81 |
| 0.5 | 35.37 | 37.37 |
D.4 Label weight distribution
The weight assignment in SPS follows the formula . In this section we make a toy experiment to explore whether such assignment is reasonable. We take KS-video as example which use top 2 classes for reconstruction. Table 14 show the results on different weight assignment. It can be seen that the performance peaks at 0.7, corresponding to the assignment in main text. Therefore the assignment of weights in the main text is justified.

### v2

@@ Appendix A Pseudo-code for MDAA
The pseudo-code of MDAA is shown in Algorithm 1. For the pre-trained models, we integrate an individual AC for each network block, using the source dataset to initialize the classifiers as well as the memory bank. During the inference and adaptation periods, the model reconstructs the output labels for each sample using the soft pseudo-label and determines which ACs need to be updated through the DLFM.
Appendix B Benchmarks,Backbone and Implementation Details
B.1. Details about the Benchmarks
@@
ALL experiments are conducted on the two popular multi-modal datasets Kinetics (Kay et al., 2017) and VGGSound (Chen et al., 2020). (Yang et al., 2024) further provides their corrupted visual and audio modality for TTA tasks.
@@
Both datasets’ visual and audio modalities were extracted following the method described in (Gong et al., 2022b). To systematically explore the distributional shifts across modalities, various corruption types were applied to both visual and audio components. Following (Yang et al., 2024), 15 corruption types were introduced for the visual modality, each with five levels of severity for comprehensive evaluation. These corruptions include “gaussian nois”, “shot noise”, “impulse noise”, “defocus blur”, “glass blur”, “motion blur”, “zoom blur”, “snow”, “frost”, “fog”, “brightness”, “contrast”, “elastic transform”, “pixelate”, and “jpeg compression”. Similarly, the audio modality was subjected to 6 different corruptions: “gaussian noise”, “traffic noise”, “crowd noise”, “rain”, “thunder”, and “wind”. The corrupted versions of these benchmarks are referred to as Kinetics50-C and VGGSound-C, respectively. Visualizations of sample corrupted video frames and audio spectrograms are provided in Figure 5 and Figure 6.
B.2. CAV-MAE Backbone
@@
CAV-MAE is employed as the pre-trained models for MM-CTTA in this paper. Its architecture consists of 11 Transformer blocks (known as feature encoder networks) dedicated to modality-specific feature extraction, followed by an additional Transformer block (known as fusion network) responsible for cross modal fusion. For the video input, 10 frames are sampled from each clip, from which a single frame is randomly selected and fed into the Transformer encoder for the visual modality. In the case of the audio input, the original 10-second audio waveform is transformed into a 2D spectrogram before being processed by the Transformer encoder for the audio modality (Gong et al., 2022a).
| KS-video | VGG-audio | |||
| 1e-3 | 2.10 | 2.08 | 0.51 | 0.47 |
| 1e-2 | 2.03 | 2.12 | 0.31 | 0.39 |
| 1e-1 | 64.87 | 69.03 | 0.32 | 0.37 |
| 1e0 | 65.43 | 69.30 | 0.36 | 0.40 |
| 1e1 | 65.27 | 69.23 | 35.82 | 35.77 |
| 1e2 | 65.29 | 68.95 | 35.78 | 35.76 |
| 1e3 | 59.32 | 59.94 | 30.36 | 30.52 |
| Method | Type | Kinetics50-C | VGGSound-C | ||||||||||||
B.3. Implementation Details
@@
In the final step of Algorithm 1, we determine the hyperparameters for MDAA. The expansion layer dimension, denoted as , theoretically benefits from larger values. However, an excessively large dimension may introduce a significant number of parameters, occupying considerable computational resources. Given the constraints of our available GPU resources, we set the dimension of to 8000. The necessity of the parameter in Eq. 3.2 has been established in (Zhuang et al., 2022). The model demonstrates stable performance over a wide range of values, indicating that as long as is within a reasonable range, its impact on the model’s performance remains minimal. Table 8 shows the results under a sweep over six orders of magnitude (i.e., ) using the same setting in the ablation study. The model performs robustly with a sufficient regularization, while collapses when the value of turns to be too extreme (larger than and smaller than ). In this paper, we set to 1 for Kinetics-50 and 10 for VGGSound. As discussed in Sections 4.2 and 4.3, the threshold in DLFM is fixed at 0.001 for both datasets, while the parameter for reconstructing soft labels is set to 2 for Kinetics-50 and 7 for VGGSound. All experiments were conducted on an RTX3090 GPU, with results averaged over three runs.
Appendix C More Experiment Results
@@
In this appendix, we provide four more experiments for reference where Appendix C.1 and C.2 are the supplementary comparative studies while Appendix C.3 is the further discussion on hyperparameters applied in MDAA.
C.1. Performance Comparison on Multi-modal Corruption
@@
We follow the experimental setup in (Yang et al., 2024) to examine whether our model remains superior to others when the adaptation task is not continuous. In this section, the corrupted target domains are independent, and the results for corrupted audio and video modalities are presented in Tables 9, 10 and 11. The results show that the SOTA performance of MDAA is due not only to its ability to combat catastrophic forgetting, but also to its strong capacity to handle MM-TTA tasks effectively.
@@
C.2. Performance Comparison on Single-modality Continual Corruption
@@
C.3. Dynamic threshold update
The threshold we used in DLFM is a fixed number. In this section we attempt to update in a dynamic way during the adaptation. We define the threshold for classifier in time as
where is the learning rate, is the initial threshold and is the batch size. is calculated to reflect the gap between Leader and follower . The original intention is to adjust the threshold size according to changes to eliminate statistical bias across domains. However, as shown in Table 15, this approach achieves lower performance while requiring more variables to be memorized. Therefore, only the fixed threshold is used in the formal method.
@@ C.3. Dynamic threshold update
Appendix D Proofs of Theorems
Proof of Theorem 1. known the optimal problem in Eqn.3.3 can be further written as:
Note above equation as , derive for as
Therefore